OpenStudy (anonymous):

evaluate limit lim x approaches infinity (1-p/x)^(qx)

OpenStudy (anonymous):

as lim x aproaches to infinity (1-p)/x =e to the power of -pq

OpenStudy (anonymous):

\[\lim x->infinity (1-p/x)^{qx}\]

OpenStudy (anonymous):

its (1-p)/x^(qx) = e^-pq

OpenStudy (anonymous):

o ok thanks

OpenStudy (anonymous):

ah, this is a good one. case 1: if q = 0 and p isn't 0. simple enough. the limit will simply be 1, because p/x -> 0 case 2: if p = 0 and q isn't 0, simple too. it will be 1 to the power of a huge number (if q is positive). if q is negative, it will be one to the power of a very small number. in either case, it will be 1. case 3: p is not zero and q is not zero make the substitution m = qx. as x ->inf, m->oo (since q > 0) thus the limit becomes (1 - pq/m)^m = e^(-pq)

OpenStudy (anonymous):

and considering when q < 0 leads to a limit of e^(-pq) too

OpenStudy (anonymous):

thats a cheap answer

OpenStudy (anonymous):

james, how do you know that the limit (1 - pq/m)^m = e^ -pq

OpenStudy (anonymous):

because (1 + x/n)^n, for very large n, converges to e^x euler's famous limit

OpenStudy (anonymous):

ok i guess you can use that result. what if you didnt know it?

OpenStudy (anonymous):

prove eulers famous limit then

OpenStudy (anonymous):

Good question. I'd use this reasoning: y = (1 - p/x)^(qx) ln(y) = qx * ln(1 - p/x) so ln(y) = ln(1 - p/x) / (1/qx) Use L'hopital's rule to calculate the limit of ln(y): = lim x->inf ( -pqx^2/(x^2 - px) ) = lim x->inf ( -pq/(1 - p/x) ) = -pq/(1 - 0) = -pq Thus ln(y) converges to -pq Since ln(y) is continous for all positive y, then y converges to e^(-pq)

OpenStudy (anonymous):

nice :)

OpenStudy (anonymous):

as for proving l'Hopital's rule, i am unsure. ask l'Hopital :)